Jefferson Lab HALL A Logbook for January 03 - All EXCEPT Automatic Entries

  • Main INDEX

    Friday

  • 94080 01/31/03 21:41 x. jiang rotate target to 180 degree
  • 94079 01/31/03 21:40 x. jiang NMR while target is at 0 degree, laser on, pt=38.08%
  • 94067 01/31/03 19:24 x. jiang left-arm DAQ , EB-1 problem, fixed by reset
  • 94048 01/31/03 16:07 Ole Hansen Shift Summary
  • 94046 01/31/03 15:58 choi Structure in dp spectra with PID cut
  • 94045 01/31/03 15:22 vasulk delta P detector cuts
  • 94030 01/31/03 13:02 jpchen target to pol. He3
  • 94023 01/31/03 11:03 Sirish compton
  • 94012 01/31/03 08:00 M Jones Owl shift summary
  • 93992 01/31/03 04:02 A. Camsonne Target : NMR Polarization 37.7%
  • 93989 01/31/03 03:34 A. Camsonne Target : moved to Pol He3
  • 93988 01/31/03 03:19 A. Camsonne Target : moved to pick-up coils position
  • 93985 01/31/03 03:12 A. Camsonne Target : pressure reference cell at the end of the run
  • 93980 01/31/03 02:47 A. Camsonne Target : Reference cell filled with N2
  • 93972 01/31/03 02:23 A. Camsonne Target : moved to empty reference cell
  • 93964 01/31/03 01:57 A. Camsonne Target : Moved to BeO
  • 93961 01/31/03 01:34 A. Camsonne Target : temperature fluctuations
  • 93960 01/31/03 01:18 A. Camsonne Target : heater temperature dropped for a few seconds
  • 93956 01/31/03 00:56 A. Camsonne Target : Holding field from 180 to 270 degrees
  • Thursday

  • 93950 01/30/03 23:58 Reitz shift summary
  • 93944 01/30/03 22:40 x. jiang NMR , polarization is 36.3%
  • 93939 01/30/03 22:05 paschke Charge Asym Sign Test results
  • 93937 01/30/03 21:52 paschke parity DAQ oversampling
  • 93931 01/30/03 21:23 paschke PZT scan
  • 93905 01/30/03 19:31 paschke pzt scans to start
  • 93897 01/30/03 17:54 kramerk Rotatable Half-Wave plate not immediately set
  • 93896 01/30/03 17:50 Reitz beam position
  • 93893 01/30/03 17:41 anonymous shift summary day, partial
  • 93879 01/30/03 16:02 R.A. Lindgre shift summary
  • 93875 01/30/03 13:23 R. Lindgren shift update
  • 93874 01/30/03 11:44 R.A. Lindgre shift update
  • 93873 01/30/03 10:22 A. Deur Sign flip check
  • 93850 01/30/03 08:00 Cusanno Shift Summary
  • 93843 01/30/03 07:29 S.Woo NMR measurement
  • Wednesday

  • 93818 01/29/03 23:58 R.A. Lindgre shift summary
  • 93817 01/29/03 23:24 x. jiang reference cell pressure change
  • 93813 01/29/03 23:03 paschke R-HRS DAQ hangup
  • 93812 01/29/03 23:00 x. jiang target direction at 270 degree
  • 93811 01/29/03 22:59 x. jiang NMR=38.45%
  • 93810 01/29/03 22:58 paschke PZT charge asymmetry scans
  • 93809 01/29/03 22:50 paschke Charge Asym Sign Test
  • 93797 01/29/03 22:16 x. jiang reference cell with N2
  • 93770 01/29/03 21:08 x. jiang target half wave plate is out.
  • 93768 01/29/03 21:07 x. jiang rotate target to 0 degree, pumping
  • 93755 01/29/03 19:04 A. Deur Plan for this evening and tonight
  • 93742 01/29/03 16:15 LeRose Shift Summary
  • 93740 01/29/03 15:47 jpchen NMR=37.9%, target rotated to 180, half wave plate moved in.
  • 93733 01/29/03 13:33 jpchen beam half wave plate in at 12:07
  • 93720 01/29/03 11:12 Sirish compton
  • 93712 01/29/03 10:20 Moffit Re: Cut Changes
  • 93711 01/29/03 09:39 jpchen NMR=39.1%, target changed from 270 to 0 degree
  • 93706 01/29/03 09:05 Moffit Parity Feedback Cut Changes
  • 93691 01/29/03 04:45 S.Woo NMR measurement
  • 93677 01/29/03 01:56 S.Woo rotated holding field and moved HWP, target
  • 93670 01/29/03 00:59 S.Woo NMR measurement
  • 93666 01/29/03 00:27 Eugene Moller results
  • Tuesday

  • 93665 01/28/03 23:55 ent shift summary
  • 93664 01/28/03 22:32 Sirish Compton
  • 93663 01/28/03 16:17 A. Deur elastic asymmetry analysis
  • 93662 01/28/03 16:07 R. Michaels tried faster clocks
  • 93654 01/28/03 15:45 A. Deur MCC will send 25 uA in Hall A
  • 93651 01/28/03 14:43 R. Michaels event builder window
  • 93648 01/28/03 11:48 vasulk 4 GeV Carbon False Asymmetry
  • 93647 01/28/03 10:25 R. Michaels modified ROC3 code
  • 93646 01/28/03 09:40 A. Deur NMR+target at zero degree
  • 93643 01/28/03 09:05 R.A. Lindgre stop run for beam studies
  • 93642 01/28/03 09:02 A. Deur Empty target in
  • 93636 01/28/03 08:00 Boeglin Owl Shift Summary
  • 93632 01/28/03 07:59 G.M. Urciuol 1/28/2003 owl shift report
  • 93600 01/28/03 00:01 ent swing shift summary
  • Monday

  • 93598 01/27/03 22:38 R. Michaels DAQ modifications
  • 93591 01/27/03 22:24 R. Michaels L-arm run numbers were stuck
  • 93586 01/27/03 22:20 Solvignon Helicity problem: all 4GeV runs are good
  • 93581 01/27/03 18:16 X. Zheng Carbon asymmetry at kine #5.3
  • 93579 01/27/03 18:01 R. Michaels re: synch problems
  • 93578 01/27/03 17:41 choi use "espace" instead of "espace_debug"
  • 93577 01/27/03 17:39 choi helicity signal synch check
  • 93574 01/27/03 15:40 Solvignon 5.3 kinematics: carbon runs asymmetry
  • 93573 01/27/03 15:37 x. jiang left arm VDC HV reset
  • 93572 01/27/03 12:22 reitz GEN timing tests
  • 93567 01/27/03 12:00 Solvignon Longitudinal lasers: wavelength adjustment
  • 93564 01/27/03 10:57 atrig Trigger-Download
  • 93563 01/27/03 10:52 atrig Trigger-Download
  • 93562 01/27/03 10:48 Solvignon Longitudinal lasers: new setting for the current
  • 93554 01/27/03 10:16 Solvignon Laser alignment
  • 93553 01/27/03 09:44 S.Woo rotated holding field to 0
  • 93552 01/27/03 08:58 S.Woo NMR measurement
  • 93543 01/27/03 08:07 Adaq Owl Shift Summary
  • 93542 01/27/03 08:07 G.M. Urciuol 1/27/03 Owl shift report
  • 93509 01/27/03 00:06 A. Deur NMR
  • Sunday

  • 93507 01/26/03 23:59 ent swing shift summary
  • 93491 01/26/03 21:25 A. Deur target spin to 180 degree
  • 93488 01/26/03 19:57 choi Synchronization problem between ROC3 and ROC10?
  • 93476 01/26/03 17:58 A. Deur NMR measurement
  • 93469 01/26/03 16:05 bogdanw shift summary
  • 93467 01/26/03 15:54 j. singh Ref Cell - N2
  • 93463 01/26/03 15:43 j. singh Ref Cell Empty run
  • 93417 01/26/03 10:50 j. singh nmr target pol. = 36.3%
  • 93402 01/26/03 08:06 Boeglin Owl shift summary
  • 93397 01/26/03 08:04 G.M.Urciuoli Target owl shift of 1/26/03
  • 93383 01/26/03 04:51 adaq Charge assymetry, 1/2 wave plate
  • Saturday

  • 93361 01/25/03 23:57 ent evening shift summary
  • 93360 01/25/03 23:57 A. Deur helicity 0 problem seems to be gone
  • 93359 01/25/03 23:40 choi Wrong user name in halog entry
  • 93358 01/25/03 23:27 Adaq Synchronization Problem
  • 93357 01/25/03 23:24 A. Deur now running in unbuffer mode on left arm
  • 93353 01/25/03 23:15 ent no ref. cell measurement
  • 93350 01/25/03 22:19 ent daq/helicity problem back
  • 93343 01/25/03 21:06 ent daq/helicity problem (left)
  • 93333 01/25/03 19:20 Solvignon Target: NMR= 41.1%
  • 93332 01/25/03 18:40 solvignon Pion rejector HV increased
  • 93329 01/25/03 17:59 solvignon Preshower/shower HV increased
  • 93326 01/25/03 16:51 Solvignon Detector Check
  • 93324 01/25/03 16:01 bogdanw shift summary
  • 93323 01/25/03 13:42 J. Singh Target NMR Pol. = 38.1
  • 93314 01/25/03 11:33 J. Singh Field Rotation 0->270
  • 93297 01/25/03 08:08 Boeglin Owl Shift Summary
  • 93293 01/25/03 07:23 vasulk NMR: 38.8%
  • 93276 01/25/03 04:56 vasulk Target Field to 0 degrees
  • 93267 01/25/03 01:45 Solvignon Preshower/shower HV increased
  • 93259 01/25/03 00:36 vasulk NMR: 40.9%
  • 93255 01/25/03 00:03 Eugene Swing shift summary
  • Friday

  • 93253 01/24/03 23:37 Solvignon Detector check
  • 93240 01/24/03 20:22 Eugene Start data taking
  • 93236 01/24/03 20:11 A. Deur Target to pol. He3 position
  • 93235 01/24/03 18:27 Eugene Beam tuning
  • 93234 01/24/03 17:41 A. Deur Target polarimetry
  • 93232 01/24/03 16:01 Ole Hansen Day Shift Summary
  • 93231 01/24/03 15:41 Moffit IHWP: OUT RHWP: 6435
  • 93230 01/24/03 14:16 Moffit wien angle = -16.2 degrees
  • 93226 01/24/03 13:42 R. Michaels Detmap record keeping
  • 93225 01/24/03 12:21 Ole Hansen Test runs
  • 93219 01/24/03 12:14 R. Michaels BPM pedestal runs -- ok
  • 93217 01/24/03 11:47 choi Right arm preshower/shower signal
  • 93208 01/24/03 11:14 R. Michaels pedrunL fixed, did pedrun 20259
  • 93195 01/24/03 08:03 Boeglin Shift Summary
  • 93191 01/24/03 06:56 vasulk NMR: 40.6%
  • 93187 01/24/03 06:31 Boeglin Right arm shower counter
  • 93154 01/24/03 00:29 Boeglin Right Arm HV Crate alarms
  • 93153 01/24/03 00:09 averett Swing Shift Summary
  • Thursday

  • 93144 01/23/03 22:32 R. Michaels ped run L-arm, new pedestals
  • 93141 01/23/03 22:18 Moffit Failed Pedestal Run: Left Arm
  • 93138 01/23/03 22:01 Moffit Another pedestal run
  • 93135 01/23/03 21:32 Moffit Right Arm: Pedestal Run
  • 93125 01/23/03 19:00 Solvignon Preshower/shower and pion rejector HV
  • 93110 01/23/03 16:46 Moffit Target to Empty Reference Cell
  • 93109 01/23/03 16:45 Moffit Target: NMR = 42.6%
  • 93105 01/23/03 16:18 A. Deur Shift summary
  • 93103 01/23/03 12:15 A. Deur 5 uA pulsed beam
  • 93102 01/23/03 11:54 jpchen correction
  • 93101 01/23/03 11:48 jpchen Target: NMR=41.2%
  • 93100 01/23/03 11:46 vasulk Rotate Field to 180
  • 93099 01/23/03 11:13 Nilanga Halo study with cuts
  • 93098 01/23/03 09:29 A. Deur beam study (35 min earlier)
  • 93091 01/23/03 08:18 A. Deur kinematics 5.2.3 done, Target spin to 180 degree
  • 93086 01/23/03 08:03 paschke shift summary
  • 93074 01/23/03 05:04 vasulk NMR: 40.9%
  • 93051 01/23/03 00:58 paschke ROC14 crash
  • Wednesday

  • 93047 01/22/03 23:56 C. C. Chang Shift Summary
  • 93040 01/22/03 23:35 choi Empty cell run analysis
  • 93030 01/22/03 21:47 C. C. Chang Beam is back.
  • 93027 01/22/03 21:44 C. C. Chang Another beam update
  • 93024 01/22/03 21:07 Wooyoung Kim NMR 43.7 %
  • 93023 01/22/03 20:09 C. C. Chang Update on beam situation
  • 93022 01/22/03 18:51 C. C. Chang Beam
  • 93021 01/22/03 17:03 Reitz bpm pedestal runs
  • 93018 01/22/03 15:57 R. Michaels day shift
  • 93017 01/22/03 15:54 LeRose Beam Status
  • 93016 01/22/03 15:24 LeRose L Q3 Nitrogen
  • 93015 01/22/03 14:39 Moffit RHWP = 6650
  • 93014 01/22/03 14:34 Moffit Beam half-wave plate IN
  • 93013 01/22/03 14:22 R. Michaels happex run 1451, no beam
  • 93010 01/22/03 12:55 LeRose beam status
  • 93009 01/22/03 12:45 LeRose beam status
  • 93006 01/22/03 12:20 R. Michaels singing wire not singing any more
  • 93003 01/22/03 12:01 Jack Left Arm VDC V plane
  • 93002 01/22/03 11:54 Moffit Target to BeO/C
  • 93001 01/22/03 11:51 R. Michaels status
  • 93000 01/22/03 10:26 Moffit Target: NMR = 40.5%
  • 92996 01/22/03 10:02 R. Michaels last prod run (for now), ref cell run
  • 92992 01/22/03 09:54 Moffit Target to Empty Ref. Cell
  • 92988 01/22/03 09:38 R. Michaels dhist correlations fixed
  • 92981 01/22/03 08:35 choi Beam halo study comments
  • 92977 01/22/03 08:25 R. Michaels re: restarting L-arm
  • 92973 01/22/03 08:11 Moffit RF trips. Beam down.
  • 92971 01/22/03 08:01 paschke Shift Summary
  • 92963 01/22/03 06:55 vasulk RTD readback with Beam on target
  • 92962 01/22/03 06:24 paschke DAQ hangs again
  • 92953 01/22/03 04:57 paschke N2 reference cell pressure (20483,1483)
  • 92952 01/22/03 04:52 paschke DAQ hang, L-HRS
  • 92946 01/22/03 04:21 vasulk NMR: 41.7% !
  • 92938 01/22/03 03:41 X. Zheng halo test - minor correction
  • 92931 01/22/03 03:29 vasulk Move to BeO target
  • 92930 01/22/03 03:27 X. Zheng Halo test
  • 92905 01/22/03 02:20 peter changed raster size
  • 92904 01/22/03 02:18 vasulk Reference cell
  • 92890 01/22/03 01:56 peter raster size and shape
  • 92883 01/22/03 01:31 paschke BeO/C target in
  • 92880 01/22/03 01:01 peter spot++ plot of beam spot for halo test.
  • 92877 01/22/03 00:51 vasulk Target to Reference cell
  • 92868 01/22/03 00:23 C. C. Chang Shift Summary
  • Tuesday

  • 92866 01/21/03 21:05 Moffit Target: Moved to BeO/C
  • 92862 01/21/03 20:39 C. C. Chang Beam spot size/halo
  • 92858 01/21/03 20:24 Moffit Target: Moved to Empty Reference Cell
  • 92857 01/21/03 19:49 Moffit Removed EPICS info from parity datafiles
  • 92854 01/21/03 18:06 Moffit Target: Moved to BeO/C
  • 92843 01/21/03 17:37 Moffit Target: Moved to Empty Reference Cell
  • 92830 01/21/03 17:14 Sirish Compton Result
  • 92829 01/21/03 17:06 Moffit Left Arm Recycled and Set
  • 92828 01/21/03 16:43 Moffit iocha14 crash. Left Arm Q2, Q3, Dipole crash
  • 92827 01/21/03 16:05 R. Michaels day shift log
  • 92825 01/21/03 15:22 Antonin Vach Changed feedback soft version
  • 92824 01/21/03 15:21 choi beam halo study
  • 92823 01/21/03 14:52 Saha Checked out Harps and 9th Dipole
  • 92822 01/21/03 14:18 R. Michaels status
  • 92821 01/21/03 14:07 RM/KP/AV Beam Studies
  • 92805 01/21/03 12:47 Solvignon EPR settings
  • 92804 01/21/03 12:37 Solvignon EPR tests with gradient coils on
  • 92801 01/21/03 12:32 R. Michaels change: 2.957
  • 92787 01/21/03 10:49 W. Kim NMR 34.86 %
  • 92786 01/21/03 10:33 R. Michaels magnetic fields changed, kine 5.2
  • 92785 01/21/03 10:16 paschke IA feedback off
  • 92784 01/21/03 10:08 R. Michaels beam studies to commence
  • 92776 01/21/03 09:01 R. Michaels scaler packaging delay fix, MCC tuning
  • 92766 01/21/03 08:25 R. Michaels beam tuning
  • 92757 01/21/03 07:58 Eugene, Pete Owl shift summary
  • 92744 01/21/03 04:38 paschke Target NMR: 34.9%
  • 92737 01/21/03 03:13 Eugene 1/2 wave plate removed
  • 92724 01/21/03 00:05 A. Deur Shif Summary
  • Monday

  • 92719 01/20/03 22:49 slifer N2 pressure for ref cell run 20456/1456
  • 92716 01/20/03 22:35 slifer target NMR: 35%
  • 92709 01/20/03 21:25 A. Deur Target polarimetry+ ref cell runs
  • 92703 01/20/03 18:29 choi espace crash resolved
  • 92702 01/20/03 18:18 A. Deur Beam really back
  • 92697 01/20/03 17:46 A. Deur beam is back
  • 92696 01/20/03 17:29 A. Deur Beam down
  • 92692 01/20/03 16:00 Ole Hansen Day Shift Summary
  • 92690 01/20/03 15:48 Ole Hansen Another replay crash
  • 92687 01/20/03 15:34 choi New espace with ring buffer decoding
  • 92686 01/20/03 15:24 Ole Hansen ESPACE crash - cannot analyze run 20450
  • 92683 01/20/03 14:34 R. Michaels parity run -- checks
  • 92665 01/20/03 12:21 Ole Hansen Beam gone
  • 92662 01/20/03 11:29 K. McCormick Target field rotated to 270 degrees
  • 92660 01/20/03 09:24 Ole Hansen Increasing beam current to 12uA
  • 92655 01/20/03 09:08 peter shift summary
  • 92653 01/20/03 07:51 Solvignon Target: NMR= 34.65%
  • 92626 01/20/03 01:34 solvignon target:NMR measurement
  • 92618 01/20/03 00:04 A. Deur Shift Summary
  • Sunday

  • 92617 01/19/03 23:16 A. Deur VDC multiplicity
  • 92608 01/19/03 20:47 Aniol pol. direct = 180.
  • 92607 01/19/03 20:38 A. Deur VDC multiplicity
  • 92606 01/19/03 20:33 vasulk Holding Field at 180 degrees
  • 92602 01/19/03 20:11 A. Deur Target spin to 180 degree. Kine 5.1.7
  • 92600 01/19/03 19:57 Antonin Vach panFFB change on parity run 1406
  • 92599 01/19/03 19:54 Antonin Vach PanFFB (minor) change
  • 92598 01/19/03 19:40 Antonin Vach RHWP scan final value
  • 92563 01/19/03 18:46 A. Deur Beam down in the Halls for orbit study
  • 92559 01/19/03 18:44 A. Deur runs 20/1435
  • 92558 01/19/03 18:25 A. Deur Beam 1/2 plate in
  • 92551 01/19/03 18:20 Antonin Vach Rotatable halfwaveplate scan
  • 92550 01/19/03 17:44 vasulk NMR: 32.6 %
  • 92549 01/19/03 17:34 A. Deur 1/2 plate change+zero crossing study
  • 92536 01/19/03 13:36 vasulk Transverse lasers (better format)
  • 92535 01/19/03 13:33 vasulk Masing procedure
  • 92534 01/19/03 13:20 vasulk Transverse Laser currents
  • 92533 01/19/03 13:09 Z.-E. Mezian Shift summary 8:00-12:00
  • 92532 01/19/03 13:07 vasulk Reboot Spectrum Analyzer
  • 92515 01/19/03 08:08 feuerbac owl shift summary
  • 92513 01/19/03 07:20 solvignon PID detector calibration: right arm
  • 92509 01/19/03 06:50 P. Solvignon NMR Measurements 35.4 %
  • 92498 01/19/03 03:22 Moffit hatsv5 crash?
  • 92492 01/19/03 02:24 Moffit Re: Enabled IA Lockserver Feedback.
  • 92488 01/19/03 01:34 Moffit/Solvi Target: NMR = 37.0% EPR = 37.9 --> 36.1%
  • 92487 01/19/03 01:16 Moffit Enabled IA Lockserver Feedback.
  • Saturday

  • 92479 01/18/03 23:53 slifer swing shift summary
  • 92472 01/18/03 23:06 slifer run number sequence
  • 92471 01/18/03 23:02 K. McCormick Reset Coda on LARM
  • 92466 01/18/03 22:40 slifer runs 20418/1418
  • 92465 01/18/03 22:18 vasulk Target RTD's Hot
  • 92458 01/18/03 19:32 W. Kim/W. Ko NMR measurement
  • 92450 01/18/03 19:04 Adaq detector replay
  • 92427 01/18/03 17:23 slifer detector replay
  • 92426 01/18/03 16:45 R. Michaels suggestion re: Qasy feedback
  • 92424 01/18/03 16:03 A. Deur Run summary
  • 92417 01/18/03 15:16 R. Michaels ET tests. Finished.
  • 92416 01/18/03 14:53 Kelleher target at pol 3He
  • 92411 01/18/03 14:30 Kelleher target to BeO
  • 92408 01/18/03 14:07 A. Deur Elastic cross section analysis
  • 92402 01/18/03 13:27 R. Michaels parity feedback problems
  • 92401 01/18/03 13:24 Kent Paschke feedback software version rolled back
  • 92400 01/18/03 13:23 Kent Paschke What to do if feedback appears to be inactive
  • 92399 01/18/03 13:22 A. Deur Right arm spectra
  • 92394 01/18/03 12:51 Kelleher target NMR = 39.49%
  • 92393 01/18/03 12:47 Kelleher Rebooted target IOC
  • 92392 01/18/03 12:38 A. Deur Beam will still be down for 30 min
  • 92391 01/18/03 12:20 Aniol happex daq restarted
  • 92389 01/18/03 11:54 A. Deur Hall B Moller
  • 92385 01/18/03 11:49 A. Deur Hall B bleed thru test
  • 92384 01/18/03 11:49 X. Zheng correction for entry
  • 92383 01/18/03 11:48 Aniol happex feedback monitor
  • 92382 01/18/03 11:47 X. Zheng Left arm PID cuts for the current kinematics (run #20410)
  • 92381 01/18/03 11:41 X. Zheng Pion rejector PID performance
  • 92380 01/18/03 11:38 X. Zheng Right arm gas cherenkov
  • 92378 01/18/03 11:07 X. Zheng Right arm PID cuts for the current kinematics (run #1406)
  • 92377 01/18/03 11:01 X. Zheng Total shower PID performance
  • 92376 01/18/03 10:50 X. Zheng Gas cherenkov PID performance
  • 92375 01/18/03 10:41 A. Deur Beam 1/2 wave plate out
  • 92365 01/18/03 09:11 Kelleher Target moved to pol 3He
  • 92359 01/18/03 08:56 Kelleher Target to BeO/C
  • 92354 01/18/03 07:46 Moffit Halog headers wrong? Target Half-wave plate
  • 92353 01/18/03 07:40 Moffit Target: NMR = 38.4%
  • 92349 01/18/03 07:31 Moffit Target to pol3he
  • 92348 01/18/03 07:30 Moffit Target: Holding field to 180, Target Halfwave plate IN
  • 92347 01/18/03 07:19 Moffit Target to pickup coils
  • 92331 01/18/03 02:58 Moffit/Kim Target: NMR = 39.1%
  • Friday

  • 92310 01/17/03 23:28 Nilanga 5 GeV 32 degree and 25 degree data
  • 92309 01/17/03 23:23 reitz start/end of run entry correctiion
  • 92298 01/17/03 22:03 Wolfgang NMR measurement
  • 92289 01/17/03 21:38 Eugene Moller results
  • 92284 01/17/03 19:28 reitz Tiefenbach vs Set Energy
  • 92280 01/17/03 19:10 Gao beam back!
  • 92279 01/17/03 18:37 vasulk NMR: 38.1 %
  • 92278 01/17/03 18:20 Reitz right Hrs
  • 92277 01/17/03 16:54 vasulk/Wolfg Gradient coils on
  • 92276 01/17/03 16:15 Higinbotham Beam Spot Size Checks
  • 92275 01/17/03 15:56 Ole Hansen Injector work
  • 92274 01/17/03 15:54 Ole Hansen More harp scans
  • 92273 01/17/03 15:51 Ole Hansen Shift Summary
  • 92271 01/17/03 14:48 Ole Hansen Ramping HRSR dipole
  • 92270 01/17/03 14:42 Jack R-arm Dynapower
  • 92269 01/17/03 14:16 choi Beam halo check
  • 92268 01/17/03 13:06 Ole Hansen Beam spot issues
  • 92267 01/17/03 12:39 Ole Hansen Harp results w raster off
  • 92251 01/17/03 12:15 Ole Hansen 12C optics data
  • 92247 01/17/03 12:14 Ole Hansen Harp scans
  • 92238 01/17/03 11:58 vasulk Target Polarization
  • 92234 01/17/03 11:38 S.Woo NMR/EPR measurement
  • 92232 01/17/03 08:04 paschke shift summary
  • 92231 01/17/03 08:04 paschke HRS-R Q1 trip
  • 92230 01/17/03 07:47 Moffit Target to BeO
  • 92229 01/17/03 07:33 paschke smoke, no fire, few hours down
  • 92228 01/17/03 07:02 paschke Fire in the Hall
  • 92221 01/17/03 05:28 paschke Feedback not working (ET server)
  • 92218 01/17/03 04:57 Moffit Target: EPR = 36.3%
  • 92217 01/17/03 04:54 Moffit Target to Pol He3
  • 92216 01/17/03 04:49 Moffit Target: NMR = 35.4%
  • 92213 01/17/03 04:27 paschke IA scan
  • 92204 01/17/03 03:50 Moffit Fill Reference Cell with N2
  • 92197 01/17/03 03:29 Moffit harpscan result
  • 92193 01/17/03 03:15 Moffit Target to Empty Reference Cell
  • 92190 01/17/03 02:47 paschke spot size
  • 92188 01/17/03 02:38 paschke harp scan
  • 92185 01/17/03 02:20 paschke Raster size
  • 92174 01/17/03 01:04 paschke tune beam on the way
  • 92173 01/17/03 00:50 Moffit Target to BeO
  • Thursday

  • 92166 01/16/03 22:22 Wolfgang Some beam in hall?
  • 92165 01/16/03 22:12 vasulk NMR: 28.9%, EPR: 30.1 %
  • 92164 01/16/03 22:05 vasulk Temperature Test Results
  • 92161 01/16/03 15:29 vasulk Run numbers
  • 92154 01/16/03 14:49 Antonin Vach feedbacks crashed at 9:30 am
  • 92152 01/16/03 14:39 vasulk Charge Asymmetry Feedback
  • 92148 01/16/03 13:51 vasulk Magnet/Spectrum Analyzer
  • 92147 01/16/03 11:56 GAO controlled access
  • 92111 01/16/03 09:54 S.Woo rotated holding field to 270
  • 92095 01/16/03 08:16 paschke Target NMR: 35.4 %, EPR: 37.1%
  • 92092 01/16/03 08:03 Eugene Chuda Owl shift summary
  • 92086 01/16/03 07:04 paschke Target field rotation to 0 degrees
  • 92077 01/16/03 04:18 paschke new feedback software version
  • 92076 01/16/03 04:09 paschke target field to 180 degrees
  • 92071 01/16/03 02:16 slifer EPR: 37.79%
  • 92070 01/16/03 02:13 paschke Target polarization (NMR) = 35.7
  • 92067 01/16/03 02:07 paschke target field rotated
  • 92066 01/16/03 02:05 paschke polarization measurement
  • 92062 01/16/03 00:53 Eugene Raster has a peak at the center
  • 92061 01/16/03 00:14 paschke/vach IA slope tuned
  • Wednesday

  • 92059 01/15/03 23:59 K. McCormick Swing Shift Summary, Jan 15
  • 92056 01/15/03 23:38 K. McCormick High Charge Asymmetry Fixed (during run 20361)
  • 92041 01/15/03 22:58 paschke IA slope has been changed
  • 92034 01/15/03 22:14 K. McCormick Right dipole reaches 600 A
  • 92033 01/15/03 22:00 K. McCormick Removed beam half-wave plate
  • 92032 01/15/03 21:58 K. McCormick Problem with Raster
  • 92009 01/15/03 19:44 P. Solvignon Polarization measurement
  • 92008 01/15/03 18:50 nanda Compton Update
  • 92007 01/15/03 16:58 K. McCormick Harp Scan with all Harps
  • 92003 01/15/03 16:37 LeRose Right Dipole Tripped Again
  • 92002 01/15/03 16:09 LeRose BTA entry
  • 92000 01/15/03 16:06 LeRose day shift summary
  • 91999 01/15/03 16:03 LeRose Harpscan
  • 91998 01/15/03 16:01 vasulk Halo Check
  • 91997 01/15/03 16:00 LeRose Harps
  • 91996 01/15/03 15:56 LeRose Harpscans
  • 91995 01/15/03 15:52 LeRose Harpscans
  • 91994 01/15/03 15:23 Solvignon EPR: polarization extraction when field drift happens
  • 91989 01/15/03 14:45 Moffit Target motion to BeO
  • 91988 01/15/03 13:36 LeRose Harp Scan Suggestion
  • 91987 01/15/03 13:06 Moffit Target: NMR = 34.3%
  • 91986 01/15/03 13:00 solvignon EPR= 35.6% and good spectrum
  • 91985 01/15/03 12:41 Moffit Target to Pick-Up Coils
  • 91984 01/15/03 11:51 LeRose Harp Scans
  • 91983 01/15/03 10:59 Solvignon EPR: more that 2% difference between down and up sweep
  • 91982 01/15/03 10:26 apar Increase Lumi HV
  • 91981 01/15/03 10:15 Moffit Holding field to 0. L-Lasers On
  • 91980 01/15/03 10:14 LeRose Harp Scan Failed
  • 91979 01/15/03 09:43 Moffit Target moved, T-lasers off
  • 91975 01/15/03 08:30 kramerk Target - NMR/EPR - 34.1/35.2%
  • 91974 01/15/03 08:04 kramerk Shift Summary
  • 91972 01/15/03 07:52 kramerk MCC is only able to deliver 2uA
  • 91971 01/15/03 07:39 paschke Polarization Measurement
  • 91950 01/15/03 04:05 kramerk Target - Rotated Field to 270
  • 91937 01/15/03 02:57 kramerk Target - NMR/EPR - 35.5/36.5
  • 91934 01/15/03 02:44 kramerk Polarization Measurement
  • 91930 01/15/03 01:20 kramerk Need access to reboot Spectral Analyzer Machine
  • Tuesday

  • 91922 01/14/03 23:53 slifer Swing shift summary
  • 91921 01/14/03 23:46 slifer run 20345 comments incorrect
  • 91914 01/14/03 22:36 solvignon target at 180deg
  • 91910 01/14/03 22:29 Solvignon Target: EPR= 38.6%, NMR= 32.9%
  • 91904 01/14/03 21:33 slifer Raster Crate
  • 91898 01/14/03 21:21 slifer Hall A parameters window/ Raster crate
  • 91896 01/14/03 20:34 Solvignon EPR with gradient coils on
  • 91889 01/14/03 19:22 nanda Compton
  • 91885 01/14/03 19:16 vasulk NMR Measurement
  • 91884 01/14/03 19:14 slifer Compton tune
  • 91880 01/14/03 18:13 slifer Beam back
  • 91873 01/14/03 18:01 slifer beam tuning
  • 91872 01/14/03 17:54 Moffit parity daq: removed v/f
  • 91871 01/14/03 17:39 apar Lumi HV ON
  • 91866 01/14/03 16:51 slifer Q2 Left arm
  • 91859 01/14/03 16:03 Kelly day shift
  • 91857 01/14/03 15:57 Kelly Q2 left trip
  • 91844 01/14/03 15:02 Moffit Junk Parity Runs 1300,1201
  • 91843 01/14/03 15:01 Eugene Moller result from Jan,09 is corrected
  • 91813 01/14/03 13:49 nanda Compton
  • 91812 01/14/03 13:20 Solvignon target at 180deg: half-wave plates at the bad angles
  • 91811 01/14/03 12:38 lkaufman Pedestal runs with HAPPEX DAQ
  • 91806 01/14/03 10:32 Solvignon laser alignment
  • 91798 01/14/03 07:56 kramerk Shift Summary
  • 91790 01/14/03 06:53 kramerk Laser Temp. Adjusted
  • 91773 01/14/03 05:26 kramerk Target - Alarm on RTD 9
  • 91766 01/14/03 04:32 kramerk QE Measurement
  • 91760 01/14/03 03:14 A. Camsonne Compton
  • 91759 01/14/03 03:10 kramerk Requested beam for Optics Studies
  • 91758 01/14/03 02:55 A.Camsonne MQA1C20
  • 91757 01/14/03 02:52 A. Camsonne Quadrupole MQA1C19
  • 91756 01/14/03 02:42 A. Camsonne MBC1C20V
  • 91755 01/14/03 02:33 A. Camsonne MBC1C20H scan
  • 91754 01/14/03 02:21 A. Camsonne MBC1C18V scan
  • 91753 01/14/03 01:31 A. Camsonne Trying to reduce background in compton chicane
  • 91752 01/14/03 00:29 kramerk Sieve Slits IN
  • 91751 01/14/03 00:20 kramerk Chaged Spectrometer Momebntum
  • 91750 01/14/03 00:04 Eugene Swing shift
  • Monday

  • 91748 01/13/03 23:48 Eugene Moller results
  • 91747 01/13/03 23:15 vasulk Holding Field Program
  • 91746 01/13/03 22:11 slifer NMR/EPR measurement (masing test)
  • 91743 01/13/03 20:28 slifer Target field rotation
  • 91738 01/13/03 17:48 slifer NMR/EPR measurement
  • 91731 01/13/03 16:24 Eugene Right arm quenched
  • 91724 01/13/03 16:05 Kelleher target field rotation
  • 91723 01/13/03 16:05 Kelly right dipole fixed!
  • 91722 01/13/03 16:02 Kelly day shift
  • 91718 01/13/03 15:47 Kelleher EPR, NMR measurements
  • 91717 01/13/03 15:45 vasulk Laser adjustments
  • 91713 01/13/03 15:11 Kelleher target at pol 3 He
  • 91706 01/13/03 14:38 Kelleher target at BeO
  • 91705 01/13/03 14:08 Moffit parity pedestal run
  • 91700 01/13/03 12:40 Moffit Parity DAQ test runs
  • 91693 01/13/03 11:09 Kelleher target movement, NMR
  • 91692 01/13/03 11:01 Moffit Parity DAQ - halog entries
  • 91690 01/13/03 10:57 apar Start_of_Run_Parity_1275
  • 91689 01/13/03 10:12 vasulk NMR polarization
  • 91681 01/13/03 07:50 Kumbartzki Owl shift Summary
  • 91668 01/13/03 04:47 Nilanga Right Arm Dipole
  • 91663 01/13/03 03:14 Nilanga Target polarization rotation
  • 91662 01/13/03 03:12 Nilanga EPR, NMR
  • 91653 01/13/03 00:05 kramerk Shift Summary
  • Sunday

  • 91651 01/12/03 23:56 kramerk Little bit of DAQ trouble
  • 91646 01/12/03 23:18 kramerk ROC14 crash
  • 91641 01/12/03 22:00 kramerk EPR Polarization - 35.7%
  • 91636 01/12/03 20:48 kramerk EPR RF left on!
  • 91633 01/12/03 20:42 nanda Compton
  • 91632 01/12/03 20:35 Kelleher target position
  • 91631 01/12/03 20:28 kramerk New Magnet Settings for 2.55 GeV
  • 91630 01/12/03 20:09 kramerk Changing Momentum to 2.55 GeV
  • 91623 01/12/03 18:09 kramerk CHL needs to make access
  • 91620 01/12/03 17:59 Kelleher target movement
  • 91615 01/12/03 17:38 kramerk Right Dipole Tripped
  • 91610 01/12/03 17:29 Kelleher target rotation motion
  • 91609 01/12/03 17:27 kramerk Controlled Access
  • 91601 01/12/03 16:02 Kelly day shift
  • 91596 01/12/03 14:40 jpchen Target NMR, moved into Pol. He3, then EPR
  • 91595 01/12/03 14:27 jpchen laser #5 had a fault light on
  • 91591 01/12/03 13:59 jpchen access trying to check transverse laser wave length
  • 91584 01/12/03 11:40 jpchen pion rejector gains
  • 91583 01/12/03 11:11 Kelly right dipole
  • 91582 01/12/03 10:46 jpchen target movde to he3 postion
  • 91577 01/12/03 10:05 jpchen target: NMR=35.59%, EPR=35.15%
  • 91574 01/12/03 09:51 jpchen reference cell filled with N2 to 147 psi
  • 91570 01/12/03 09:29 jpchen target back to 0 for NMR, EPR then to reference (empty)
  • 91564 01/12/03 07:58 R. Michaels owl shift summary
  • 91559 01/12/03 06:51 R. Michaels halog scaler oddities fixed
  • 91554 01/12/03 06:05 R. Michaels R-arm dipole trips again
  • 91553 01/12/03 05:45 A. Camsonne Target : spectral analyser
  • 91549 01/12/03 05:27 A. Camsonne Target : Holding Field to 270 degrees
  • 91547 01/12/03 05:10 R. Michaels scaler checkout
  • 91546 01/12/03 04:38 R. Michaels R-arm dipole update
  • 91537 01/12/03 02:21 A. Camsonne Target Moved back to Pol He3
  • 91534 01/12/03 02:10 R. Michaels R-arm dipole trip
  • 91533 01/12/03 01:32 nanda Compton Update
  • 91532 01/12/03 01:31 R. Michaels finished kinematics 4.4, go to 4.5
  • Saturday

  • 91527 01/11/03 23:57 A. Deur Shift summary
  • 91526 01/11/03 23:55 A. Deur How to reset Q2
  • 91523 01/11/03 23:41 A. Deur Q2 left reset
  • 91522 01/11/03 23:32 Kelleher polarization
  • 91521 01/11/03 23:04 A. Deur Q2 left trip
  • 91516 01/11/03 22:14 Kelleher field rotation
  • 91512 01/11/03 20:48 Kelleher target rotation
  • 91508 01/11/03 19:59 Kelleher target position
  • 91504 01/11/03 19:13 Kelleher reference cell
  • 91499 01/11/03 18:32 A. Deur helicity signal
  • 91496 01/11/03 18:04 Kelleher NMR, EPR
  • 91491 01/11/03 17:22 Moffit/Vache IA Scan
  • 91488 01/11/03 17:05 A. Deur lot of helicity 0 events
  • 91485 01/11/03 16:06 glashausser Shift Summary 1/11/03 day
  • 91483 01/11/03 16:04 solvignon target: EPR=29.9%
  • 91479 01/11/03 15:50 Kelly replay for setting 4.3.3 para
  • 91477 01/11/03 15:14 Kelly OTR 1C12
  • 91472 01/11/03 14:01 solvignon target at 270deg
  • 91463 01/11/03 12:06 solvignon target: pol. 3He position
  • 91462 01/11/03 12:05 solvignon target: BeO position
  • 91461 01/11/03 12:04 solvignon target: NMR= 31%, EPR= 32.1%
  • 91456 01/11/03 10:41 solvignon what we should do when we want to run the target at 180deg
  • 91455 01/11/03 10:33 solvignon target: rotate at 0deg
  • 91454 01/11/03 10:28 solvignon target: NMR= 32.5%, EPR= 29.2%
  • 91453 01/11/03 10:27 solvignon target: go to 180deg
  • 91452 01/11/03 10:12 Adaq beam half wave plate positions
  • 91447 01/11/03 09:30 Higinbotham Harp 5 & 6 Problems (follow-up from Friday evening)
  • 91444 01/11/03 09:06 solvignon Target: NMR= 32.6%, EPR= 34.8%
  • 91439 01/11/03 08:08 A. Camsonne Target : filling reference cell N2
  • 91434 01/11/03 07:59 R. Michaels owl shift summary
  • 91432 01/11/03 07:44 R. Michaels A_q feedback converged
  • 91431 01/11/03 07:35 A. Camsonne Target : Reference cell (empty) at 10 uA
  • 91428 01/11/03 07:27 A. Camsonne Target Moved to reference cell empty
  • 91418 01/11/03 06:48 R. Michaels IA slope too small, spot size too big
  • 91417 01/11/03 06:30 A. Camsonne Target : on BeO since 5:54
  • 91416 01/11/03 06:18 Markowitz rotatable wave plate vs bcm1 asy, half wave plate out
  • 91415 01/11/03 06:08 Markowitz rotation of wave plate vs. Q asy
  • 91414 01/11/03 06:06 A. Camsonne Target : Reference cell RTD 9 at 56 degrees
  • 91413 01/11/03 06:03 R. Michaels set the IA slope, enable feedback
  • 91412 01/11/03 05:56 R. Michaels No need to re-do IA, proceed with BeO
  • 91409 01/11/03 05:48 R. Michaels Mistaken RWP value used
  • 91406 01/11/03 05:45 R. Michaels IA scan
  • 91399 01/11/03 05:31 A. Camsonne Target : reference cell RTD 9 at 55
  • 91392 01/11/03 05:21 R. Michaels putting waveplate in, G0 optimum set
  • 91391 01/11/03 05:19 Markowitz Putting in the 1/2 wave plate
  • 91389 01/11/03 05:10 R. Michaels RWP scan
  • 91388 01/11/03 05:07 R. Michaels HAPPEX vs Spectrometers
  • 91387 01/11/03 04:57 A. Camsonne Target : Reference cell RTD 9 at 52 degrees
  • 91383 01/11/03 04:36 R. Michaels correction: we ARE in chicane
  • 91381 01/11/03 04:32 A. Camsonne Target : Reference cell RTD 9
  • 91380 01/11/03 04:30 A. Camsonne Target : Polarization EPR : 32% NMR 29.9 %
  • 91376 01/11/03 04:17 R. Michaels beam not in chicane
  • 91372 01/11/03 03:57 Markowitz half wave plate being removed
  • 91369 01/11/03 03:52 A. Camsonne Target : RTD9 still increasing
  • 91368 01/11/03 03:48 R. Michaels Spectrometer and HAPPEX agreement
  • 91361 01/11/03 03:25 R. Michaels switched root output to adaql1 disk
  • 91358 01/11/03 03:14 A. Camsonne Target alarm : RTD 9 reference cell alarm
  • 91356 01/11/03 03:12 R. Michaels BCM2 pedestals
  • 91349 01/11/03 02:40 A. Camsonne Target alarm : RTD 9 reference cell alarm
  • 91344 01/11/03 02:00 R. Michaels warmup runs
  • 91343 01/11/03 01:52 A. Deur sign check runs with new espace
  • 91338 01/11/03 00:49 R. Michaels switch parity disks, status
  • 91337 01/11/03 00:28 Higinbotham Shift Summary
  • 91335 01/11/03 00:09 choi New ESPACE helicity
  • Friday

  • 91334 01/10/03 23:39 Higinbotham Harp Scans from Last Night
  • 91333 01/10/03 22:36 vasulk Field direction check
  • 91332 01/10/03 22:10 vasulk Polarized Helium-3 Alarms
  • 91331 01/10/03 21:44 Solvignon Work on target
  • 91330 01/10/03 19:11 nanda Compton
  • 91329 01/10/03 18:55 kramerk Spectral Analyzer Update
  • 91328 01/10/03 17:15 paschke Unser monitor signal to Parity DAQ
  • 91327 01/10/03 17:15 vasulk Pol. Helium-3 position
  • 91326 01/10/03 17:10 slifer Target gradient coils
  • 91325 01/10/03 16:32 Eugene, Xiao Pion rejector
  • 91323 01/10/03 16:16 slifer Pion Rejector
  • 91322 01/10/03 15:31 K. McCormick RTD 6 Readback
  • 91317 01/10/03 15:08 kramerk Spectral Analyzer unstable
  • 91316 01/10/03 14:04 paschke IA system waveplate changed
  • 91315 01/10/03 12:02 Ole Hansen Computer work
  • 91311 01/10/03 08:02 R. Michaels owl shift summary
  • 91310 01/10/03 07:49 A. Camsonne Target : moved to reference cell ( empty )
  • 91309 01/10/03 07:48 A. Camsonne Target polarization : NMR 34.24% EPR 35.6%
  • 91308 01/10/03 07:03 Adaq Target : moving to pick-up coils and changed holding field to 0
  • 91305 01/10/03 06:42 A. Camsonne Target : turned laser 2 and 3 on
  • 91304 01/10/03 06:32 A. Camsonne Target : Polarization NMR 34.8% EPR 36.9%
  • 91303 01/10/03 06:27 A. Camsonne Target : RTD readout at 270 degrees
  • 91300 01/10/03 06:17 A. Camsonne Target : turned laser 2 and 4 on
  • 91299 01/10/03 06:08 A. Camsonne Target : moved to Pol He3 and changing holding field to 270 degrees
  • 91297 01/10/03 05:55 A. Camsonne Target : moving to pick-up coil
  • 91296 01/10/03 05:53 R. Michaels Qasy feedback works.
  • 91295 01/10/03 05:53 A. Camsonne Target : Pressure fluctuation
  • 91290 01/10/03 05:20 A. Camsonne Target : Filling reference cell with N2 at 147 psi
  • 91289 01/10/03 05:09 R. Michaels Reference cell running
  • 91284 01/10/03 04:41 A. Camsonne Target : Moving to reference cell (empty)
  • 91278 01/10/03 03:37 R. Michaels re : Qasy fdbk stoppage
  • 91277 01/10/03 03:32 feuerbac Missing pion rejector channel
  • 91276 01/10/03 03:30 R. Michaels Qasy feedback working again
  • 91275 01/10/03 03:20 R. Michaels beam waveplate moved in
  • 91274 01/10/03 03:19 feuerbac hot VDC channel
  • 91269 01/10/03 03:01 R. Michaels Qasy feedback stopped
  • 91268 01/10/03 02:15 A. Deur Beautiful He3 pressure curve
  • 91267 01/10/03 02:09 R. Michaels IA feedback adjustments
  • 91266 01/10/03 01:51 K. Slifer Target polarization EPR : 35.3 %
  • 91265 01/10/03 01:44 A. Camsonne Target Polarization NMR : 33.3%
  • 91262 01/10/03 01:21 R. Michaels Download new scaler code
  • 91261 01/10/03 01:09 R. Michaels R-arm settled, stop run *248
  • 91257 01/10/03 00:41 R. Michaels enabled parity A_q feedback
  • 91256 01/10/03 00:35 R. Michaels kinematics 4.2.3
  • 91255 01/10/03 00:32 Eugene Moller target masking incident
  • 91252 01/10/03 00:28 A. Deur Ref cell orientation
  • 91251 01/10/03 00:25 A. Camsonne Target moved from BeO to Pol He3
  • 91249 01/10/03 00:22 Eugene Problem with Moller helicity delay solved
  • 91247 01/10/03 00:17 Eugene Moller results
  • 91244 01/10/03 00:03 vasulk NMR: 30.5%, EPR: 31.8%
  • 91243 01/10/03 00:02 Higinbotham Shift Summary
  • Thursday

  • 91242 01/09/03 23:55 A. Deur BCM down and up difference
  • 91241 01/09/03 23:01 Antonin Vach IA slope calibration
  • 91240 01/09/03 22:48 Eugene G0 influence
  • 91239 01/09/03 21:31 choi EPICS reading at the end of run
  • 91232 01/09/03 21:01 Higinbotham Raster Check
  • 91231 01/09/03 20:52 Higinbotham Harp5 Not Working
  • 91225 01/09/03 20:12 Antonin Vach RHWP scans with hall C
  • 91224 01/09/03 18:17 vasulk NMR, EPR polarization
  • 91223 01/09/03 18:08 vasulk Laser # 7 current set to 15 A
  • 91221 01/09/03 16:50 paschke Rotating Waveplate setpoint (source)
  • 91220 01/09/03 16:01 Saha Run Summary
  • 91217 01/09/03 14:55 gilman added warning
  • 91216 01/09/03 14:43 gilman helicity setup
  • 91213 01/09/03 14:29 Lingyan target rotated to 0 degree
  • 91208 01/09/03 13:35 R. Michaels scaler cabling change (DATA FORMAT CHANGE)
  • 91205 01/09/03 13:12 Lingyan target: transverse laser realignment
  • 91203 01/09/03 12:45 vasulk R-arm and L-arm Reactz
  • 91202 01/09/03 10:47 Moffit Half-Wave Plate Scan
  • 91199 01/09/03 10:05 Lingyan target encoder positions
  • 91190 01/09/03 09:24 Lingyan target polarization and rotation
  • 91184 01/09/03 08:59 A. Camsonne Owl shift summary
  • 91180 01/09/03 08:02 Solvignon RTD#4 dead
  • 91129 01/09/03 04:12 A. Camsonne Spot ++ Raster at target
  • 91126 01/09/03 04:04 A. Camsonne Spectrometer momentum 1.051 Gev/c
  • 91125 01/09/03 03:56 Solvignon Target: EPR=23%
  • 91121 01/09/03 02:57 A. Deur Beam position
  • 91118 01/09/03 02:49 Solvignon Target in transverse: problem
  • 91110 01/09/03 02:04 A. Camsonne Spot at target after injector change
  • 91107 01/09/03 01:37 Solvignon Target: NMR 20.9%, EPR: 13.93%
  • 91106 01/09/03 00:49 A. Deur N2 pressure curve analysis
  • 91104 01/09/03 00:02 Reitz shift summary
  • Wednesday

  • 91087 01/08/03 22:54 A. Deur Ref Cell N2 contamination
  • 91079 01/08/03 21:51 Reitz Q2 on left arm tripped
  • 91065 01/08/03 21:18 Nilanga NMR, EPR, go to 180
  • 91061 01/08/03 21:03 Reitz correction start/end of run
  • 91058 01/08/03 20:59 Higinboham Beam Thru Compton Chicane
  • 91054 01/08/03 19:06 paschke/kauf IA system re-calibration
  • 91035 01/08/03 18:07 R. Michaels bad TDC 1875, signif. data loss
  • 91025 01/08/03 17:30 paschke/kauf IA system re-calibration
  • 91022 01/08/03 17:10 A. Deur peak at y_tgt=0
  • 91021 01/08/03 16:46 R. Michaels sign check result
  • 91018 01/08/03 16:32 choi G0 helicity direct reporting for a short period
  • 91013 01/08/03 16:22 vasulk NMR: 39.23 %, EPR: 39.93%
  • 91012 01/08/03 16:20 paschke/kauf IA waveplate scan
  • 91001 01/08/03 16:02 LeRose Shift Summary
  • 90978 01/08/03 13:33 LeRose controlled access
  • 90977 01/08/03 12:15 vasulk Pol. Helium-3 position
  • 90976 01/08/03 11:57 vasulk NMR: 40.2 %, EPR: 38.7%
  • 90963 01/08/03 09:41 Lingyan raster change
  • 90931 01/08/03 08:02 kramerk Shift Summary
  • 90927 01/08/03 07:45 kramerk Beam Half Wave Plate is IN
  • 90925 01/08/03 07:42 solvignon detector check with quasielastic data
  • 90916 01/08/03 07:16 A. Deur Double scattering contamination
  • 90902 01/08/03 06:22 solvignon Target polarization measurement
  • 90896 01/08/03 06:02 solvignon target: in reference cell position
  • 90895 01/08/03 06:01 kramerk New Magnet Settings
  • 90894 01/08/03 05:54 kramerk Requested Beam from MCC
  • 90893 01/08/03 05:39 solvignon target in pick-up coils position
  • 90892 01/08/03 05:28 kramerk Changing the spectrometer settings
  • 90891 01/08/03 05:20 kramerk Asked for Controlled Access
  • 90887 01/08/03 05:05 kramerk Reference Cell filled with 37 psig he3
  • 90883 01/08/03 04:54 kramerk New Magnet Settings
  • 90879 01/08/03 04:27 kramerk Changed Magnet Settings
  • 90875 01/08/03 04:18 kramerk New Spectrometer Settings
  • 90874 01/08/03 03:20 kramerk Evacuating Ref Cell
  • 90873 01/08/03 03:19 kramerk Small Leak in N2 Refcell
  • 90872 01/08/03 03:17 kramerk Change Spectrometer Magnet
  • 90868 01/08/03 03:07 solvignon target IOC crashed
  • 90858 01/08/03 02:18 solvignon target in ref. cell position
  • 90857 01/08/03 02:04 kramerk Requested a new beam position for refcell
  • 90856 01/08/03 01:56 kramerk New Spectrometer Settings
  • 90850 01/08/03 01:31 Higinbotham Bull's Eye Harp Data
  • 90849 01/08/03 01:29 kramerkChn Changing the Spectrometer Momentum
  • 90848 01/08/03 01:27 Higinbotham Bull's Eye Scan
  • Tuesday

  • 90797 01/07/03 23:50 Higinbotham Hall Access (Target DAQ)
  • 90792 01/07/03 23:41 Higinbotham Helicity "Sign Check" Data
  • 90791 01/07/03 23:32 moffit NMR 37.29% EPR 41.35%
  • 90780 01/07/03 21:53 Eugene Moller results
  • 90777 01/07/03 21:09 vasulk Target IOC Network connection
  • 90776 01/07/03 17:23 Eugene /adaqs3/data2 died
  • 90775 01/07/03 17:13 vasulk Temperature Test Results
  • 90774 01/07/03 16:02 Saha Day Shift Summary
  • 90772 01/07/03 14:52 R. Michaels DAQ problems all fixed !
  • 90771 01/07/03 14:47 gilman Q1 trip reasons
  • 90766 01/07/03 11:43 paschke IA feedback system not working
  • 90765 01/07/03 11:25 paschke,kauf IA slope measurement
  • 90764 01/07/03 11:10 Lingyan target thermometer RTD6 up
  • 90741 01/07/03 08:34 solvignon target: polarization measurement
  • 90732 01/07/03 08:06 kramerk Real Shift Summary
  • 90728 01/07/03 07:51 kramerk MCC needs to reboot IOC
  • 90727 01/07/03 07:41 kramerk Stopped Beam to do Polarization Measurement
  • 90726 01/07/03 07:28 kramerk Moved Sieve Slit in
  • 90711 01/07/03 06:10 R. Michaels R-arm NOT dead (but no evt scalers)
  • 90708 01/07/03 06:01 solvignon Target analysis: EPR-NMR calibration
  • 90707 01/07/03 05:01 kramerk Change of Program
  • 90706 01/07/03 04:53 R. Michaels R-arm DAQ is dead
  • 90696 01/07/03 04:16 solvignon Target: move to BeO
  • 90689 01/07/03 03:57 kramerk Slight Difficulty adjusting Q2 momentum
  • 90683 01/07/03 03:31 kramerk Right Arm Q2 tripped
  • 90682 01/07/03 03:13 kramerk Spectrum Analyzer Not Working
  • 90681 01/07/03 03:06 solvignon target: RTD
  • 90680 01/07/03 02:52 solvignon target: reference cell RTD
  • 90679 01/07/03 02:50 solvignon Target: .RTD check
  • 90674 01/07/03 02:08 kramerk Controlled Access Update
  • 90666 01/07/03 01:25 kramerk Shift Summary
  • 90653 01/07/03 00:37 R. Michaels ESPACE helicity and DAQ
  • 90651 01/07/03 00:06 Higinbotham Shift Summary
  • Monday

  • 90650 01/06/03 23:58 vasulk Temperature Test
  • 90649 01/06/03 23:18 choi No problem with helicity signal
  • 90648 01/06/03 23:14 Higinbotham Beam Spot without Raster
  • 90647 01/06/03 22:11 R. Michaels DAQ update.
  • 90646 01/06/03 22:04 choi Bad helicity signal
  • 90637 01/06/03 20:50 solvignon Target: EPR in longitudinal is fine
  • 90630 01/06/03 20:03 Higinbotham First Harp Scan (spot too big)
  • 90624 01/06/03 19:43 averett Q2 right trip
  • 90623 01/06/03 19:29 vasulk RTD 6 fluctuation
  • 90622 01/06/03 19:14 vasulk Target moved to BeO position
  • 90621 01/06/03 18:54 vasulk Target Encoder positions
  • 90618 01/06/03 17:45 R. Michaels R-arm DAQ problems
  • 90606 01/06/03 17:27 vasulk Reference cell RTD
  • 90596 01/06/03 16:06 jpchen day shift summary
  • 90579 01/06/03 14:25 jpchen In beam permit
  • 90578 01/06/03 14:11 R. Michaels large deadtimes last night
  • 90577 01/06/03 12:59 jpchen BCM offset problem fixed
  • 90576 01/06/03 12:32 K. W. Q1 tripped
  • 90575 01/06/03 12:27 krishniw Magnet alarms
  • 90574 01/06/03 11:54 Lingyan Harp Scan(continued)
  • 90571 01/06/03 11:44 jpchen Q1 magnet
  • 90570 01/06/03 11:44 Lingyan Harp Scan
  • 90568 01/06/03 11:34 jpchen controlled access ended
  • 90567 01/06/03 11:34 R. Michaels TIR resistors removed
  • 90564 01/06/03 11:32 choilef left arm V2 plane
  • 90557 01/06/03 11:12 Lingyan magnet reset
  • 90556 01/06/03 11:11 jpchen longitudinal EPR works now
  • 90555 01/06/03 11:00 jpchen laser power circuit break reset
  • 90550 01/06/03 09:56 R. Michaels run 20077 (L-arm) DAQ test
  • 90548 01/06/03 09:22 R. Michaels Left arm DAQ problems
  • 90544 01/06/03 09:02 moffit G0 Rotatable Half-Wave Plate Scan
  • 90542 01/06/03 08:48 Lingyan laser cameras
  • 90526 01/06/03 03:36 solvignon target: BeO + carbon position
  • 90522 01/06/03 01:26 Solvignon EPR in transverse
  • 90521 01/06/03 00:46 vasulk New Target Encoder Positons
  • 90520 01/06/03 00:07 choi magnet trips
  • Sunday

  • 90516 01/05/03 21:11 vasulk Hall in Controlled Access
  • 90515 01/05/03 19:57 choi espace slow write fixed
  • 90514 01/05/03 19:00 vasulk Magnets set
  • 90513 01/05/03 18:34 vasulk He3 bottle Open
  • 90512 01/05/03 17:46 jpchen end of controlled access
  • 90511 01/05/03 17:42 jpchen left Q1 adn dipole
  • 90510 01/05/03 17:40 jpchen reference cell and cooling jet
  • 90509 01/05/03 14:37 kramerk Hall A in Controlled Access
  • 90491 01/05/03 13:03 kramerk MCC sending CW beam
  • 90490 01/05/03 12:53 kramerk MCC sending tune beam
  • 90489 01/05/03 10:54 choi espace spectra/write too slow
  • 90488 01/05/03 10:35 choi Detectors, DAQ, espace.
  • 90487 01/05/03 09:23 kramerk NMR - AFP RF Coil Problems
  • Saturday

  • 90484 01/04/03 15:18 R. Michaels timing change (final, I think)
  • 90471 01/04/03 12:22 R. Michaels timing change, preshower problem
  • 90470 01/04/03 12:17 atrig Trigger-Download
  • Friday

  • 90457 01/03/03 12:18 R. Michaels QRT fixed
  • 90450 01/03/03 11:46 R. Michaels helicity QRT signal bad
  • Thursday

  • 90436 01/02/03 12:23 R. Michaels scaler and document updates
  • Wednesday